OpenStudy (anonymous):

What type of property is this? (9*-4) * 7 = 7 * (9*-4)

OpenStudy (anonymous):

its equal to itself

OpenStudy (anonymous):

Have a paper that says commutative, associate, identity, inverse, and distributive.

OpenStudy (anonymous):

i think identity then

OpenStudy (mathstudent55):

Notice that inside the parentheses, the 9 * (-4) remained the same. Think of 9 * (-4) as y. What property does y * 7 = 7 * y show?

OpenStudy (anonymous):

because if its equal to itself then its identical

OpenStudy (anonymous):

well if u divide both sides by 7 then u have y = y

OpenStudy (anonymous):

Commutative multiplication?

OpenStudy (mathstudent55):

All that changed was the order of the multiplication. Commutative: you can change order Associative: you can change grouping Identity: a * 1 = a (it's a property of 1 in multiplication). Inverse: a * 1/a = 1 (reciprocals) Distributive: a(b + c) = ab + ac

OpenStudy (mathstudent55):

Correct. To me it looks like the commutative property of multiplication.

OpenStudy (anonymous):

Thank you. That's what I had. Just wanted to double check.

OpenStudy (mathstudent55):

You're welcome.